G-Animal’s Journal

1983, Vol. 15, No. 4

Fanstord University

Contents

Aamport The Gnats and Gnus . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Oz, Yannakakis An Sorting Algorithm . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

Revisions 5

Larry Manmaker The Definitive Computer Manual (Masterly) . . . . . . . . . . . . . . . 5

The Gnats and Gnus

A Document Preparation System^{∗}

Leslie A. Aamport^{†}

Keywords: Gnats, Gnus

Abstract Starting from a historical point of view …

1 Why Gnu

According to [?]^{1} …

• Some items are – bullets.

1. And some items (a) are numbers.

• But some items are gnus A sort of large aninals.

1.1 Everything is changing

Change 1: Even fonts are changing.

∗to Mary

†to Peter

1 Note

\underline{Aamport, 3}

The figure is typeset in \figurefont. The caption body is typeset in

\captionbodyfont. The header of the caption is typeset in

                   \captionheaderfont.

Fig. 1: The body of the caption.

An Sorting Algorithm^{∗}

Wizard V. Oz^{†} Mihalis Yannakakis^{‡}

Keywords: Bubble, Sort

Abstract

The joy of sorting …

1 O(n log n/log log n)

See^{1}

2 Conclusions

[?] …

∗to bubble. †Fanstord University ‡University of Provence 1 Some gnu!

Revisions

Larry Manmaker The Definitive Computer Manual. Golden edition, 1993, Chips-R-Us: Silicon Valley.

The golden oldies … Édouard Masterly Stanford University